# Beyond Grids: Multi-objective Bayesian Optimization With Adaptive Discretization

### v3

@@ 4.5 Termination
If ${\cal S}_{t}=\emptyset$ at the beginning of round $t$, then the algorithm terminates. We show in the appendix that at the latest, the algorithm terminates when $\overline{\omega}_{t}\leq\min_{j}\epsilon^{j}$. Upon termination, it returns a non-empty set $\hat{\mathcal{P}}$ of decided nodes together with the corresponding $\bm{\epsilon}$-accurate Pareto set $\hat{P}=\bigcup_{x_{h,i}\in\hat{\mathcal{P}}}X_{h,i}~{},$ which is the union of the cells corresponding to the nodes in $\hat{\mathcal{P}}$.
5 Sample complexity bounds
Assumption 1.
Theorem 1.
Let $\bm{\epsilon}=[\epsilon^{1},\ldots,\epsilon^{m}]^{\mathsf{T}}$ be given with $\epsilon=\min_{j\in[m]}\epsilon^{j}>0$. Let $\delta\in(0,1)$ and $\bar{D}>D_{1}$. For each $h\geq 0$, let $V_{h}\in\tilde{O}(\rho^{\alpha h})$; for each $\tau\in\mathbb{N}$, let $\beta_{\tau}\in O(\log(\tau^{2}/\delta))$; the exact definitions are given in the appendix. When we run Adaptive $\bm{\epsilon}$-PAL with prior $GP(0,\bm{k})$ and noise $\mathcal{N}(0,\sigma^{2})$, the following holds with probability at least $1-\delta$.
$\min$$\left\{K_{1}\beta_{T}T^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log T\right)^{\frac{-(\bar{D}+\alpha)}{\bar{D}+2\alpha}}+K_{2}T^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log T\right)^{\frac{\alpha}{\bar{D}+2\alpha}},\sqrt{\frac{C\beta_{T}\gamma_{T}}{T}}\right\}\leq\epsilon~{},$
Sketch of the proof of Theorem 1: We divide the proof of Theorem 1 into three essential parts. First, we prove that the algorithm terminates in finite time, and examine the events that necessitate and the ones that follow termination. Here, it is important to note that if the largest uncertainty diameter in a given round $t$ is less than or equal to $\epsilon$, then the algorithm terminates. This introduces a way on how to proceed on upper bounding sample complexity, namely, we can upper bound the sum of these uncertainty diameters over all rounds. Then, by observing that i) the term $V_{h}$ decays to 00 as $h$ grows indefinitely, which means that the algorithm refines up until a finite number of tree levels, and that ii) a node cannot be refined more than a finite number of times before expansion, we can conclude that the algorithm terminates in finite time. After this point we prove, using Hoeffding bounds, that the true function values live inside the uncertainty hyper-rectangles with high probability. We then proceed on first proving the dimension-type sample complexity bounds and then the information-type bounds. We use the aforementioned termination condition and upper bound the sum of the uncertainty diameters over all rounds. We use Cauchy-Schwarz inequality to manipulate the expression as we desire and then upper bound it in terms of information gain using an already established result in the paper. For the dimension-type bounds, we consider expressing the sum over rounds as a sum over levels $h$ of the tree of partitions and upper bound it using the notion of metric dimension. We take the minimum of the two bounds, thus achieving the bound stated in Theorem 1.
Remark 2.
Note that we minimize over two different bounds in Theorem 1. The term that involves $\gamma_{T}$ corresponds to the information-type bound, while the other term corresponds to the metric dimension-type bound. Equivalently, we can express our information-type bound as $\tilde{O}(g(\epsilon))$ where $g(\epsilon)=\min\{T\geq 1:\sqrt{\gamma_{T}/T}\leq\epsilon\}$ and our metric dimension-type bound as $\tilde{O}(\epsilon^{-(\frac{\bar{D}}{\alpha}+2)})$ for any $\bar{D}>D_{1}$. For certain kernels, such as squared exponential and Matérn kernels, $\gamma_{T}$ can be upper bounded by a sublinear function of $T$ (see Srinivas et al. (2012)). Our information type-bound is of the same form as in Zuluaga et al. (2016). When ${\cal X}$ is a finite subset of the Euclidean space, we have $D_{1}=0$, and thus, our metric-dimension type bound becomes near-$O(1/\epsilon^{2})$, which is along the same lines with the almost optimal, gap-dependent near-$O(1/\text{gap}^{2})$ bound for Pareto front identification in Auer et al. (2016).
Proposition 1.
6 Computational complexity analysis
Discarding. This phase can be further separated into two substeps. In the first substep, the pessimistic Pareto front is determined by choosing the leaf nodes whose lower confidence bounds are not dominated by any other point in the leaf set. In the second phase, the leaf points whose upper confidence bounds are $\bm{\epsilon}$-dominated by pessimistic Pareto set points are discarded. If done naively by comparing each point with all the other points, both of the phases can result in $\mathcal{O}(n^{2})$ complexity. However, there are efficient methods that achieve lower computational complexity. In our implementation, we adopt an efficient algorithm in Kung et al. (1975) to reduce computational complexity to $\mathcal{O}(n\log n)$ when $m=2$ or $m=3$. Another algorithm proposed by Kung et al. (1975) can be implemented and it achieves $\mathcal{O}(n(\log n)^{m-2}+n\log n)$ complexity for $m>3$.
7 Experiments
This section empirically evaluates Adaptive $\epsilon$-PAL, comparing its performance and efficiency against state-of-the-art multi-objective Bayesian optimization (MOBO) methods. We focus on validating the effectiveness of the adaptive discretization strategy and assessing the algorithm’s ability to find an $\bm{\epsilon}$-accurate Pareto set sample-efficiently.
7.1 Performance metrics
We use a combination of three performance metrics: $\epsilon$-accuracy ratio, $\epsilon$-coverage ratio, and average mean-squared error (MSE). Since the objective functions tested are continuous, the true Pareto front contains infinitely many points. To compute the metrics, we approximate the true Pareto front by sampling $10,000$ points uniformly from the input space, evaluating the objective function at these points, and identifying the Pareto optimal designs among them. We ensure the sampling density is high (gap between consecutive samples less than $\epsilon/10$) so that this discretized front serves as a reasonable ground truth.
$\bm{\epsilon}$-coverage ratio. This metric measures how well the predicted set $\hat{P}$ covers the true Pareto front. It is computed as the ratio of true Pareto points (from the discretized approximation) that are $\bm{\epsilon}$-dominated by at least one point in $\bm{f}(\hat{P})$ to the total number of true Pareto points. Equivalently, we check if a true Pareto point is within $2\bm{\epsilon}$ (in the sense of $\preceq_{2\bm{\epsilon}}$) of the closest predicted point.
Average mean-squared error (MSE). This metric complements $\bm{\epsilon}$-coverage by quantifying the average closeness of the predicted front to the true front. It is computed by averaging the squared Euclidean distance between each true Pareto point (in the objective space) and the closest predicted Pareto point in $\bm{f}(\hat{P})$. Ideally, we seek high values for both accuracy and coverage, and a low value for MSE.
7.2 Multi-objective Bayesian optimization algorithms
We compare Adaptive $\bm{\epsilon}$-PAL with three state-of-the-art MOBO methods: PESMO (Hernández-Lobato et al., 2016), ParEGO (Knowles, 2006), and USeMO (Belakaria et al., 2020). Since PESMO, ParEGO, and USeMO require a pre-specified evaluation budget, while Adaptive $\bm{\epsilon}$-PAL terminates based on its confidence criteria, we set the budget for the competitors to be at least as large as the number of evaluations performed by Adaptive $\bm{\epsilon}$-PAL upon termination in our experiments. This ensures competitors have access to at least as much information. We used the implementations for PESMO and ParEGO from the Spearmint Bayesian optimization library [^3] https://github.com/HIPS/Spearmint/tree/PESM and for USeMO from the authors’ open-source code [^4] https://github.com/belakaria/USeMO.
7.3 Simulation setup & results
Results. We evaluate the predicted Pareto front points returned by each algorithm using the metrics defined above. Table 2 shows the average of the $\bm{\epsilon}$-accuracy and $\bm{\epsilon}$-coverage ratios, calculated for different evaluation thresholds $\bm{\epsilon}^{\prime}$. Using smaller $\bm{\epsilon}^{\prime}$ values provides a stricter assessment.
$\bm{\epsilon}^{\prime}$ Adaptive $\bm{\epsilon}$-PAL ($h_{\max}=24$) Adaptive $\bm{\epsilon}$-PAL ($h_{\max}=10$) Adaptive $\bm{\epsilon}$-PAL ($h_{\max}=9$) PESMO ParEGO USeMO
0.050 $98$ p m 2 $99$ p m 1 $99$ p m 1 $92$ p m 2 $94$ p m 1 $92$ p m 2
0.010 $97$ p m 2 $98$ p m 2 $97$ p m 1 $90$ p m 3 $94$ p m 1 $74$ p m 4
0.005 $97$ p m 1 $97$ p m 1 $90$ p m 3 $90$ p m 3 $84$ p m 4 $60$ p m 5
0.001 $78$ p m 4 $64$ p m 5 $42$ p m 5 $54$ p m 5 $56$ p m 5 $26$ p m 4
Table 3: Average mean-squared error (MSE) and total running time (hh:mm:ss) of the algorithms. Lowest MSE and runtime are bolded.
Metric Adaptive $\bm{\epsilon}$-PAL ($h_{\max}=24$) Adaptive $\bm{\epsilon}$-PAL ($h_{\max}=10$) Adaptive $\bm{\epsilon}$-PAL ($h_{\max}=9$) PESMO ParEGO USeMO
MSE (×$\times 10^{-6}$) 5 8 40 20 30 4500
Runtime 15:15:24 00:01:00 00:00:27 00:09:20 00:07:30 00:23:40
Comparing with competitors, Adaptive $\bm{\epsilon}$-PAL (even with $h_{\text{max}}=10$) achieves superior combined accuracy and coverage across the different evaluation thresholds. USeMO shows good accuracy for $\epsilon^{\prime}=0.05$ but degrades quickly and exhibits poor coverage, likely due to returning a sparse set of points. PESMO and ParEGO offer reasonable performance, with ParEGO slightly better overall in this setup, but neither consistently matches Adaptive $\bm{\epsilon}$-PAL’s performance, especially at stricter thresholds.
Table 3 presents the average MSE and total running time. The MSE results corroborate the accuracy/coverage findings: Adaptive $\bm{\epsilon}$-PAL achieves the lowest MSE, indicating its predicted front is closest to the true front on average. The runs with $h_{\text{max}}=10$ and $h_{\text{max}}=24$ yield comparable MSE values, significantly better than competitors. USeMO has a notably high MSE, consistent with its poor coverage.
8 Conclusion
Appendix
@@ 8 Conclusion
Algorithm 1 Adaptive $\bm{\epsilon}$-PAL
1:$\mathcal{X}$, $(\mathcal{X}_{h})_{h\geq 0}$, $(V_{h})_{h\geq 0}$, $\bm{\epsilon}$, $\delta$, $(\beta_{\tau})_{\tau\geq 1}$; GP prior $\bm{\mu}_{0}=\bm{\mu}$, $\bm{k}_{0}=\bm{k}$
@@
Appendix A The proof of Theorem 1
@@
The proof of Theorem 1 is composed of a series of sophisticated steps, and is divided into multiple subsections. First, we describe in Section A.1 a set of preliminary results that will be utilized in obtaining dimension-type and information-type bounds on the sample complexity. Then, in Section A.2, we prove a key result that provides a sufficient condition for the termination of Adaptive $\bm{\epsilon}$-PAL. We also show in this section that Adaptive $\bm{\epsilon}$-PAL returns an $\bm{\epsilon}$-accurate Pareto set when it terminates and bounds the maximum depth node that can be created by the algorithm before it terminates. In the proof, $\tau_{s}$ represents the number of evaluations performed by the algorithm until termination and $t_{s}$ represents the round (iteration) at which the algorithm terminates. Throughout the proof, for a given $\bm{\epsilon}$, we let $\epsilon=\min_{j}\epsilon^{j}$, and assume that $\bm{\epsilon}$ is such that $\epsilon>0$. Unless noted otherwise, all inequalities that involve random variables hold with probability one.
A.1 Preliminary results

### v4

@@ 4.5 Termination
If ${\cal S}_{t}=\emptyset$ at the beginning of round $t$, then the algorithm terminates. We show in the appendix that at the latest, the algorithm terminates when $\overline{\omega}_{t}<\min_{j}\epsilon^{j}$. Upon termination, it returns a non-empty set $\hat{\mathcal{P}}$ of decided nodes together with the corresponding $\bm{\epsilon}$-accurate Pareto set $\hat{P}=\bigcup_{x_{h,i}\in\hat{\mathcal{P}}}X_{h,i}~,$ which is the union of the cells corresponding to the nodes in $\hat{\mathcal{P}}$. The full procedure is given in Algorithm 1.
Algorithm 1 Adaptive $\bm{\epsilon}$-PAL
@@
5 Sample complexity bounds
Assumption 1.
Theorem 1.
Let $\bm{\epsilon}=[\epsilon^{1},\ldots,\epsilon^{m}]^{\mathsf{T}}$ be given with $\epsilon=\min_{j\in[m]}\epsilon^{j}>0$. Let $\delta\in(0,1)$ and $\bar{D}>D_{1}$. For each $h\geq 0$, let
$V_{h}=4C_{\bm{k}}(v_{1}\rho^{h})^{\alpha}\left(\sqrt{C_{2}+2\log\left(\frac{2h^{2}\pi^{2}m}{6\delta}\right)+h\log N+\left(\frac{-4D_{1}}{\alpha}\log\left(C_{\bm{k}}(v_{1}\rho^{h})^{\alpha}\right)\right)^{+}}+C_{3}\right)~,$
for some strictly positive constants $C_{2}$ and $C_{3}$, where $x^{+}:=\max\{0,x\}$ for $x\in\mathbb{R}$. Moreover, for each $\tau\in\mathbb{N}$, define $\beta_{\tau}=2\log(2m\pi^{2}N^{h_{max}+1}{{(\tau+1)}^{2}}/(3\delta))$. When we run Adaptive $\bm{\epsilon}$-PAL with prior $GP(0,\bm{k})$ and noise $\mathcal{N}(0,\sigma^{2})$, the following holds with probability at least $1-\delta$:
$\min$ $\left\{K_{1}\beta_{T}T^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log T\right)^{\frac{-(\bar{D}+\alpha)}{\bar{D}+2\alpha}}+K_{2}T^{\frac{-\alpha}{\bar{D}+2\alpha}}\left(\log T\right)^{\frac{\alpha}{\bar{D}+2\alpha}},\sqrt{\frac{C\beta_{T}\gamma_{T}}{T}}\right\}<\epsilon~,$
Sketch of the proof of Theorem 1: We divide the proof of Theorem 1 into three essential parts. First, we prove that the algorithm terminates in finite time, and examine the events that necessitate and the ones that follow termination. Here, it is important to note that if the largest uncertainty diameter in a given round $t$ is less than or equal to $\epsilon$, then the algorithm terminates. This introduces a way on how to proceed on upper bounding sample complexity, namely, we can upper bound the sum of these uncertainty diameters over all rounds. Then, by observing that i) the term $V_{h}$ decays to 0 as $h$ grows indefinitely, which means that the algorithm refines up until a finite number of tree levels, and that ii) a node cannot be refined more than a finite number of times before expansion, we can conclude that the algorithm terminates in finite time. After this point we prove, using Hoeffding bounds, that the true function values live inside the uncertainty hyper-rectangles with high probability. We then proceed on first proving the dimension-type sample complexity bounds and then the information-type bounds. We use the aforementioned termination condition and upper bound the sum of the uncertainty diameters over all rounds. We use Cauchy-Schwarz inequality to manipulate the expression as we desire and then upper bound it in terms of information gain using an already established result in the paper. For the dimension-type bounds, we consider expressing the sum over rounds as a sum over levels $h$ of the tree of partitions and upper bound it using the notion of metric dimension. We take the minimum of the two bounds, thus achieving the bound stated in Theorem 1.
Remark 2.
Note that we minimize over two different bounds in Theorem 1. The term that involves $\gamma_{T}$ corresponds to the information-type bound, while the other term corresponds to the metric dimension-type bound. Equivalently, we can express our information-type bound as $\tilde{O}(g(\epsilon))$ where $g(\epsilon)=\min\{T\geq 1:\sqrt{\gamma_{T}/T}<\epsilon\}$ and our metric dimension-type bound as $\tilde{O}(\epsilon^{-(\frac{\bar{D}}{\alpha}+2)})$ for any $\bar{D}>D_{1}$. For certain kernels, such as squared exponential and Matérn kernels, $\gamma_{T}$ can be upper bounded by a sublinear function of $T$ (see Srinivas et al. (2012)). Our information type-bound is of the same form as in Zuluaga et al. (2016). When $\cal X$} is a finite subset of the Euclidean space, we have $D_{1}=0$, and thus, our metric-dimension type bound becomes near-$O(1/\epsilon^{2})$, which is along the same lines with the almost optimal, gap-dependent near-$O(1/\text{gap}^{2})$ bound for Pareto front identification in Auer et al. (2016).
Proposition 1.
6 Computational complexity analysis
Discarding. This phase can be further separated into two substeps. In the first substep, the pessimistic Pareto front is determined by choosing the leaf nodes whose lower confidence bounds are not dominated by any other point in the leaf set. In the second phase, the leaf points whose upper confidence bounds are $\bm{\epsilon}$-dominated by pessimistic Pareto set points are discarded. If done naively by comparing each point with all the other points, both of the phases can result in $\mathcal{O}(n^{2})$ complexity. However, there are efficient methods that achieve lower computational complexity. In our implementation, we adopt Algorithm 3.1 of Kung et al. (1975) to reduce computational complexity to $\mathcal{O}(n\log n)$ when $m=2$ or $m=3$. For $m>3$, one can use Algorithm 4.1 of Kung et al. (1975) to achieve a $\mathcal{O}(n(\log n)^{m-2})$ complexity.
7 Experiments
This section empirically evaluates Adaptive $\epsilon$-PAL, comparing its performance and efficiency against other multi-objective Bayesian optimization (MOBO) methods. We focus on validating the effectiveness of the adaptive discretization strategy and assessing the algorithm’s ability to find an $\bm{\epsilon}$-accurate Pareto set sample-efficiently.
7.1 Performance metrics
We use a combination of three performance metrics: $\epsilon$-accuracy ratio, $\epsilon$-coverage ratio, and average mean-squared error (MSE). Since the objective functions tested are continuous, the true Pareto front contains infinitely many points. To compute the metrics, we approximate the true Pareto front by sampling $10,000$ points uniformly from the input space, evaluating the objective function at these points, and identifying the Pareto optimal designs among them.
[…]
7.2 Multi-objective Bayesian optimization algorithms
We compare Adaptive -PAL with four state-of-the-art MOBO methods: PESMO (Hernández-Lobato et al., 2016), ParEGO (Knowles, 2006), USeMO (Belakaria et al., 2020), and qNEHVI (Daulton et al., 2021). The latter is a hypervolume-based method specifically designed to handle noisy observations by integrating over the posterior uncertainty of the true Pareto front. Since PESMO, ParEGO, USeMO, and qNEHVI require a pre-specified evaluation budget, while Adaptive -PAL terminates based on its confidence criteria, we set the budget for the competitors to be at least as large as the number of evaluations performed by Adaptive -PAL upon termination in our experiments. This ensures competitors have access to at least as much information. We used the implementations for PESMO and ParEGO from the Spearmint Bayesian optimization library [^4] https://github.com/HIPS/Spearmint/tree/PESM, for USeMO from the authors’ open-source code [^5] https://github.com/belakaria/USeMO, and for qNEHVI from the BoTorch library. [^6] https://botorch.org/docs/multi_objective
7.3 Simulation setup & results
qNEHVI. We run the Noisy Expected Hypervolume Improvement (qNEHVI) algorithm for 100 iterations. We use the sequential setting (batch size $q=1$) and the official implementation in BoTorch. The acquisition function is optimized using multi-start L-BFGS-B, consistent with the other baselines.
Results. We evaluate the predicted Pareto front points returned by each algorithm using the metrics defined above. Table˜2 shows the average of the $\bm{\epsilon}$-accuracy and $\bm{\epsilon}$-coverage ratios, calculated for different evaluation thresholds $\bm{\epsilon}^{\prime}$. Using smaller $\bm{\epsilon}^{\prime}$ values provides a stricter assessment.
$\bm{\epsilon}^{\prime}$
Adaptive $\bm{\epsilon}$-PAL
($h_{\max}=24$)
Adaptive $\bm{\epsilon}$-PAL
($h_{\max}=10$)
Adaptive $\bm{\epsilon}$-PAL
($h_{\max}=9$)
PESMO ParEGO USeMO    qNEHVI
0.050 $98$ p m 2 $99$ p m 1 $99$ p m 1 $92$ p m 2 $94$ p m 1 $92$ p m 2 $99$ p m 1
0.010 $97$ p m 2 $98$ p m 2 $97$ p m 1 $90$ p m 3 $94$ p m 1 $74$ p m 4 $96$ p m 1
0.005 $97$ p m 1 $97$ p m 1 $90$ p m 3 $90$ p m 3 $84$ p m 4 $60$ p m 5 $95$ p m 2
0.001 $78$ p m 4 $64$ p m 5 $42$ p m 5 $54$ p m 5 $56$ p m 5 $26$ p m 4 $60$ p m 4
Table 3: Average mean-squared error (MSE) and total running time ($h$:$m$:$s$) of the algorithms. Lowest MSE and runtime are bolded.
Metric
Adaptive $\bm{\epsilon}$-PAL
($h_{\max}=24$)
Adaptive $\bm{\epsilon}$-PAL
($h_{\max}=10$)
Adaptive $\bm{\epsilon}$-PAL
($h_{\max}=9$)
PESMO ParEGO USeMO qNEHVI
MSE ($\times 10^{-6}$) 5 8 40 20 30 4500 10
Runtime 15:15:24 00:01:$0$ 00:00:27 00:09:20 00:07:30 00:23:40 00:00:40
Comparing with competitors, Adaptive $\bm{\epsilon}$-PAL (even with $h_{\text{max}}=10$) achieves superior combined accuracy and coverage across most evaluation thresholds. qNEHVI matches the best performance at $\epsilon^{\prime}=0.05$ and significantly outperforms all other baselines at stricter thresholds. Nonetheless, Adaptive $\bm{\epsilon}$-PAL with its theoretical parameters ($h_{\text{max}}=24$) still achieves the highest accuracy at the most stringent thresholds of $\epsilon^{\prime}\leq 0.005$. USeMO shows good accuracy for $\epsilon^{\prime}=0.05$ but degrades quickly, likely due to returning a sparse set of points. PESMO and ParEGO offer reasonable performance but are consistently outperformed by both qNEHVI and Adaptive $\bm{\epsilon}$-PAL.
Table˜3 presents the average MSE and total running time. The MSE results corroborate the accuracy/coverage findings: Adaptive $\bm{\epsilon}$-PAL with $h_{\text{max}}=24$ achieves the lowest MSE, indicating its predicted front is closest to the true front on average. The qNEHVI and Adaptive $\bm{\epsilon}$-PAL ($h_{\text{max}}=10$) runs yield the next-best MSE values, significantly outperforming other competitors. USeMO has a notably high MSE, consistent with its poor coverage. Observing the running times reveals the practical benefit of tuning $h_{\text{max}}$. Reducing $h_{\text{max}}$ from $24$ to $10$ or $9$ decreased the runtime dramatically while preserving strong performance. qNEHVI is also highly efficient, with a runtime of just 40 seconds, faster than all baselines except the highly-tuned Adaptive $\bm{\epsilon}$-PAL ($h_{\text{max}}=9$). This demonstrates that Adaptive $\bm{\epsilon}$-PAL can be made exceptionally fast by selecting a practical tree depth limit, while still offering an accuracy-advantage over state-of-the-art methods.
8 Conclusion
Acknowledgments
This work was supported by the Scientific and Technological Research Council of Türkiye (TÜBİTAK) under Grant 215E342. The work of C. Tekin was supported by the BAGEP Award of the Science Academy; by the Turkish Academy of Sciences Distinguished Young Scientist Award Program (TÜBA-GEBİP-2023); by TÜBİTAK 2024 Incentive Award.
Appendix
Appendix A Structure and dimensionality of the design space
Before we initiate our theoretical analysis, we first provide some technical definitions to be used throughout. We begin by defining well-behaved metric spaces. Our results will hold under any well-behaved metric space.
Definition 10.
  1. 1.
  2. 2.
  3. 3.
    We assume that the cells have geometrically decaying radii, i.e., there exist $0<\rho<1$ and $0<v_{2}\leq 1\leq v_{1}$ such that we have $B(x_{h,i},\,v_{2}\rho^{h})\subseteq X_{h,i}\subseteq B(x_{h,i},\,v_{1}\rho^{h})$ for every $h\geq 0$. Note that we have $2v_{2}\rho^{h}\leq\textup{diam}(X_{h,i})\leq 2v_{1}\rho^{h}$, where $\textup{diam}(X_{h,i})=\textup{sup}_{x,y\in X_{h,i}}d(x,y)$.
The first property implies that, for every $h\geq 0$, the cells $X_{h,i},\,1\leq i\leq N^{h}$ partition $\mathcal{X}$. This can be observed trivially by reductio ad absurdum. The second property intuitively means that, as $h$ grows, we get a more refined partition. The third property implies that the nodes $x_{h,i}$ are evenly spread out in the space.
Definition 11.
Appendix B The proof of Theorem 1
@@
The proof of Theorem 1 is composed of a series of sophisticated steps, and is divided into multiple subsections. First, we describe in Section B.1 a set of preliminary results that will be utilized in obtaining dimension-type and information-type bounds on the sample complexity. Then, in Section B.2, we prove a key result that provides a sufficient condition for the termination of Adaptive $\bm{\epsilon}$-PAL. We also show in this section that Adaptive $\bm{\epsilon}$-PAL returns an $\bm{\epsilon}$-accurate Pareto set when it terminates and bounds the maximum depth node that can be created by the algorithm before it terminates. In the proof, $\tau_{s}$ represents the number of evaluations performed by the algorithm until termination and $t_{s}$ represents the round (iteration) at which the algorithm terminates. Throughout the proof, for a given $\bm{\epsilon}$, we let $\epsilon=\min_{j}\epsilon^{j}$, and assume that $\bm{\epsilon}$ is such that $\epsilon>0$. Unless noted otherwise, all inequalities that involve random variables hold with probability one.
B.1 Preliminary results